### v1

@@ 3.3 Data Analysis
Figure 4: Each dot represents the standard deviation of the qMT parameter for one participant. The mean standard deviation was calculated across the three reconstruction methods. Linear regression was performed to analyze the increase of the standard deviation with increasing motion. Black stars denote slopes that differ significantly from zero (* for p-value < 0.05, ** for p-value < 0.01). The slopes of this linear regression analysis are further analyzed in Fig. 5.
@@
Figure 5: Analysis of the parameter estimates’ standard deviation in various ROIs as a function of the motion score (cf. Fig. 4). The histogram shows the slope of a linear regression, normalized by the respective intercept. This analysis pools all (6) qMT parameters and 4 ROIs. The proposed contrast-optimized basis results in smaller slopes compared to the SVD basis, indicating better motion correction.
4 Results
Fig. 2 shows estimates of two representative motion parameters for two participants. The scan on the left is virtually motion-free with translations well below 1 mm and rotations below 1°. The motion score of this case was 0.5 mm when using the contrast-optimized basis. In this example, the motion estimates based on the contrast-optimized basis appear less noisy compared to the SVD basis. The corresponding parameter maps, particularly the $R_{x}$ map (Fig. 3), and the $R_{1}^{f}$ and $R_{1}^{s}$ maps (Supporting Fig. S3) exhibit slightly reduced artifacts and enhanced details, which suggests that the contrast-optimized basis allows for accurate motion estimates. Assuming that motion is negligible in this case, we can use the reconstruction without motion correction as a reference and we observe visually a better agreement between the motion correction with the proposed basis and the reference as compared to the SVD basis (cf. Fig. 3).
In the second exemplary case - referred to as the “moving” participant - the motion score was 1.66 mm when using the proposed basis, which corresponds to the $87^{th}$ percentile of all 86 participants. The motion parameters estimated using the contrast-optimized basis appear also less noisy compared to the SVD basis. Further, we observe substantial systematic deviations (cf. right column of Fig. 2 and Supporting Fig. S2 for all motion parameters). In this case, the contrast-optimized basis elicits a substantially improved image quality in the quantitative maps, as exemplified in the $m_{0}^{s}$ maps in Fig. 3. The remaining qMT parameter maps for both participants can be found in Supporting Figs. S3 and S4.
To evaluate the performance of the motion correction across all 85 scans in our dataset, we calculated the standard deviation for the above-mentioned ROIs and analyzed them as a function of the respective pair-wise motion score (Fig. 4). Without motion correction, the standard deviation consistently increases with increasing motion, and the slope of a linear regression model differs from zero at a significance level of 0.01 for the majority parameters and ROIs (see also Supporting Fig. S5). When performing motion correction, the motion-induced parameter variability is reduced substantially, with most of the regression slopes being non-significantly different from zero, which indicates effective motion correction. Visually, the proposed basis outperforms the SVD basis in most ROIs and parameters. To confirm this improvement, we normalized the slope of each parameter and ROI by the respective intercept (no motion) and compared the three reconstructions, pooled over all qMT parameters (6) and ROIs (4) (Fig. 5). This analysis also suggests that the proposed basis outperforms the SVD basis, which is confirmed by a paired t-test, which revealed a significant reduction in the normalized slope ($p<0.01$) when using the proposed basis instead of the SVD basis.
5 Discussion
We proposed to enhance the contrast-to-noise ratio between brain parenchyma and CSF by rotating the SVD basis. To this end, we used a generalized eigendecomposition, which is inspired by Region-Optimized Virtual Coils (ROVir).9 We demonstrated that the increased contrast improves the motion estimates compared to an SVD basis, leading to better quality parameter maps.
@@ 5 Discussion
The proposed basis can be used as a one-to-one replacement for a traditional SVD basis. Therefore, both approaches have the same target applications, foremost transient-state quantitative MRI techniques, such as MR-Fingerprinting, where the same spin dynamics are repeated while filling the k-space. Further, the proposed approach comes at no additional computation costs during the reconstruction.
@@
We demonstrated that the contrast-optimized basis significantly improves the motion parameter estimation. Although we tested our method only on brain images and limited our investigation to rigid motion correction, the contrast-optimized basis can also be created for other body parts with two distinct tissue types. Like with the original SVD approach, key factors for successful implementation include 3D imaging to mitigate through-slice motion and a k-space trajectory that provides adequate coverage in each repetition of the spin dynamics, facilitating a time-segmented reconstruction. In this study, we used a koosh-ball trajectory with golden-angle increments, which repeatedly samples the center of k-space, and paired it with a TV-regularized low-resolution reconstruction to mitigate undersampling artifacts.
@@
While the proposed motion correction substantially reduced the artifacts in the parameter maps, Figs. 3 suggests that, in cases of severe motion, the image quality is still impaired despite motion correction with either basis. This is more evident in Supporting Fig. S6, where the quantitative maps are severely degraded by the strong motion artifacts, which resulted in a motion score of 6.35 mm (motion parameters shown in Supporting Fig. S7): despite the considerable improvements obtained after motion correction (especially in the $R_{1}^{f}$ maps), the quantitative maps are still severely affected by motion artifacts. One explanation might be the inherently low temporal resolution of self-navigated motion correction. In our current implementation, each motion state is assigned every 4 s block, and improving the temporal resolution will be part of future work.
@@
The low temporal resolution entails the assumption that no motion occurs during a 4 s block. To address this limitation, we discard data before and after strong motion. However, in cases of continuous motion, the trade-off between motion and undersampling artifacts imposes a ceiling on the motion correction performance.
@@ 6 Conclusions
We propose a contrast-optimized basis function for self-navigated motion correction in quantitative MR. We utilize the generalized eigendecomposition to increase the contrast-to-noise ratio between brain tissues and CSF, which improves the accuracy of the motion estimates and, ultimately, the image quality of the quantitative parameter maps. The proposed technique does not require any sequence modifications and/or additional scan time. Consequently, it can be seamlessly integrated into various quantitative MRI methods, e.g., inversion recovery or multi-echo spin echo, where signal variations over time can be effectively captured in a low-rank subspace.
@@
Financial Disclosure
The authors have no financial disclosure to report.
@@
The reconstruction pipeline described in Section 3.3 is implemented in Julia v1.10, except for the volume registration (as part of the motion estimation) which is performed using SPM in MATLAB. The codes to generate the proposed contrast-optimized basis are available on GitHub.
@@
Figure S1: Percentage of data removal for all the datasets estimated using a motion score6 on the motion estimates derived using the SVD basis7 and the proposed contrast-optimized basis.
Figure S2: Sagittal view of the remaining qMT parameters maps ($R_{1}^{f}$, $R_{2}^{f}$, $R_{1}^{s}$, and $T_{2}^{s}$) of a scan with virtually no motion (“still” participant), with the three reconstruction methods showing visually identical results.
Figure S3: Motion parameters estimated using the SVD basis7 and the proposed contrast-optimized basis in case of no voluntary motion (“still” participant) and in presence of voluntary motion (“moving” participant). The former led to no apparent degradation of the parametric maps, while the latter produced motion artifacts which are mitigated by the motion correction methods (Fig. 3 and Supporting Fig. S4).
Figure S4: Sagittal view of the remaining qMT parameters maps ($R_{1}^{f}$, $R_{2}^{f}$, $R_{1}^{s}$, and $T_{2}^{s}$) of a representative acquisition, where the subject exhibited a motion score at the $87^{\text{th}}$ percentile (“moving” participant), with the three reconstruction methods showing visually identical results.
@@
Figure S5: Each dot represents the standard deviation of the qMT parameter for one participant. The mean standard deviation was calculated across the three reconstruction methods. Linear regression was performed to analyze the increase of the standard deviation with increasing motion. Black stars denote slopes that differ significantly from zero (* for p-value < 0.05, ** for p-value < 0.01). The motion correction using the estimates derived using the SVD basis provide a considerable reduction in the slopes of the linear regression curves, with most being non-significantly different from zero, which indicates effective motion artifact compensation. The slopes of this linear regression analysis are further analyzed in Fig. 5.
Figure S6: Sagittal view of qMT parameters maps in presence of strong motion caused by voluntary movement of the participant (Supporting Fig. S7). Motion correction with the original SVD basis substantially reduced motion artifacts, which are further reduced when using the proposed contrast-optimized basis.
Figure S7: Motion parameters estimated using the SVD basis7 and the proposed contrast-optimized basis in case of mild motion (“moving” participant), and in presence of large motion (motion score of 6.35 mm and 5.85 mm for the proposed contrast-optimized basis and the SVD basis respectively), which led to substantial degradation of the parametric maps, partially mitigated by the motion correction methods (Supporting Fig. S6).
Supporting Information
Appendix S1 Consensus-based motion estimates
The motion estimates relative to each sequence were repeatedly estimated, where each volume served as a reference to co-register all the remaining volumes (29) per each of the six flip angle patterns. This iterative approach was preferred over selecting a single reference per each sequence, to mitigate the risk of the chosen volume being affected by motion artifacts.
A weighted consensus estimate was then computed to minimize the influence of outliers. For a set of motion parameters $\{p_{i}\}$, the consensus estimate $p_{\text{cons}}$ is iteratively updated as:
$p_{\text{cons}}^{(k+1)}=\frac{\sum_{i}w_{i}^{(k)}p_{i}}{\sum_{i}w_{i}^{(k)}},$
with $k$ being the number of iterations, and the weights $w_{i}$ are inversely proportional to the residual distances, $r_{i}=\|p_{i}-p_{\text{cons}}^{(k)}\|$, given by:
$w_{i}^{(k)}=\frac{1}{1+r_{i}}.$
The iterative process continues until the changes in $p_{\text{cons}}$ fall below a tolerance threshold. The so-derived affine matrices are then used for the k-space-based motion correction.
Additionally, to further improve the accuracy of our motion estimates, we co-registered high-resolution volumes derived from each of the six flip angle patterns. To this extent, we combined all the data from each flip angle pattern (2 min of acquisition), and reconstructed it in the subspace of the proposed contrast-optimized basis using the OptISTA algorithm 25 and locally low-rank regularization 29, 30, 31 to reduce residual noise on the final image. The resulting 1 mm isotropic volumes (6 in total, one per each sequence) were then co-registered using mri_robust_register from Freesurfer,32 with the second sequence volume as reference. The motion estimates were then applied to the individual 4 s blocks of each flip angle pattern.
data removed (%)
Frequency
Figure S1: Percentage of data removal for all the datasets estimated using a motion score 21 on the motion estimates derived using the SVD basis 8 and the proposed contrast-optimized basis.
Figure S2: Motion parameters estimated using the SVD basis 8 and the proposed contrast-optimized basis in case of no voluntary motion (“still” participant) and in presence of voluntary motion (“moving” participant). The former led to no apparent degradation of the parametric maps, while the latter produced motion artifacts which are mitigated by the motion correction methods (Fig. 3 and Supporting Fig. S4).
Figure S3: Sagittal view of the remaining qMT parameters maps ($R_{1}^{f}$, $R_{2}^{f}$, $R_{1}^{s}$, and $T_{2}^{s}$) of a scan with virtually no motion (“still” participant), with the three reconstruction methods showing visually identical results.
Figure S4: Sagittal view of the remaining qMT parameters maps ($R_{1}^{f}$, $R_{2}^{f}$, $R_{1}^{s}$, and $T_{2}^{s}$) of a representative acquisition, where the subject exhibited a motion score at the $87^{\text{th}}$ percentile (“moving” participant), with the three reconstruction methods showing visually identical results.
Figure S5: Each dot represents the standard deviation of the qMT parameter for one participant. The mean standard deviation was calculated across the three reconstruction methods. Linear regression was performed to analyze the increase of the standard deviation with increasing motion. Black stars denote slopes that differ significantly from zero (* for p-value < 0.05, ** for p-value < 0.01). The motion correction using the estimates derived using the SVD basis provide a considerable reduction in the slopes of the linear regression curves, with most being non-significantly different from zero, which indicates effective motion artifact compensation. The slopes of this linear regression analysis are further analyzed in Fig. 4.
Figure S6: Sagittal view of qMT parameters maps in presence of strong motion caused by voluntary movement of the participant (Supporting Fig. S7). Motion correction with the original SVD basis substantially reduced motion artifacts, which are further reduced when using the proposed contrast-optimized basis.
Figure S7: Motion parameters estimated using the SVD basis 8 and the proposed contrast-optimized basis in case of mild motion (“moving” participant), and in presence of large motion (motion score of 6.35 mm and 5.85 mm for the proposed contrast-optimized basis and the SVD basis respectively), which led to substantial degradation of the parametric maps, partially mitigated by the motion correction methods (Supporting Fig. S6).

### v2

@@ 3.4 ROI Analysis
Figure 6: Analysis of noise-like artifacts across all 85 datasets. Each dot represents the standard deviation of the qMT parameter for one participant. The standard deviation was calculated for each motion correction version. Linear regression was performed to analyze the increase in the standard deviation with increasing motion. Black stars denote slopes that differ significantly from zero (* for p-value < 0.05, ** for p-value < 0.01). The slopes of this linear regression analysis are further analyzed in Fig. 7.
4 Results
4.1 Simulated Motion
Fig. 2 compares motion estimates, using the SVD and the proposed contrast-optimized basis, to the ground truth. Due to the limited temporal resolution, both motion estimates fail to capture high-frequency components of the ground truth motion. Nonetheless, they approximate the slow components of the motion well. By visually comparing the estimates based on the two bases, we observe a better approximation when using the proposed basis, particularly during the rapid changes of motion states (cf. magnifications in Fig. 2). This improvement is corroborated by a lower RMSE score (cf. Eq. 13) when using the proposed basis (0.35 mm) compared to the SVD basis (0.38 mm). We note that the RMSE score captures the rapid oscillations of the ground truth motion, providing a lower bound of 0.3 mm caused by the estimates’ temporal resolution of 4 s. This bound was calculated by downsampling the ground truth. An additional case is shown in Supplementary Fig. S6-S7, where the RMSE score decreased from 1.33 mm (SVD) to 1.26 mm (proposed). For the other two motion patterns, the RMSE scores are 0.49/0.48 mm and 0.75/0.66 mm (SVD/proposed).
Motion correction with either estimates resulted in a significant recovery of image quality, as shown in the $m_{0}^{s}$ and $R_{x}$ maps in Fig. 3. When using the proposed approach as compared to the SVD basis, the improved accuracy of the motion estimates translates to enhanced image quality in the parametric maps (cf. magnifications in Fig. 3). The improvements are most apparent in the $R_{x}$ map (g and h vs. e). Still, the $m_{0}^{s}$ map also exhibits a subtle bias to lower values in the gray matter when using the SVD basis (c vs. a) that is reduced when using the proposed approach (d).
All motion parameters and the remaining parametric maps are reported in Supplementary Fig. S4-S5. Further, an example with stronger simulated motion (motion score of 6.39 mm) is provided in Supplementary Figs. S6-S7, which resulted in an improvement of the RMSE score from 1.33 mm to 1.26 mm using the proposed basis.
4.2 Inherent Motion
Fig. 4 shows estimates of representative motion parameters for two participants. The scan on the left has minimal motion, with translations well below 1 mm and rotations below 1°. The motion score of this case was 0.41 mm when using the contrast-optimized basis, which is the smallest across our 85 acquisitions (0th percentile motion score). The corresponding parametric maps (Fig. 5), including the $R_{1}^{f}$ and $R_{1}^{s}$ maps (Supporting Fig. S9) exhibit slightly reduced artifacts and enhanced details, which suggests that the contrast-optimized basis enables accurate motion estimates. In particular, the $R_{x}$ map (Fig. 5) shows fewer hyperintense regions and appears nearly isointense between gray and white matter, which is the expected contrast 18 (cf. Fig. 3a, e).
The second exemplary case has strong motion (motion score 10.53 mm, corresponding to the 100th percentile, calculated with the proposed basis) and we observe substantial artifacts when reconstructing without motion correction (cf. right column of Fig. 4 and Supporting Fig. S8 for all motion parameters). Motion correction with either basis substantially improves image quality. Comparing the maps reconstructed with either basis, we, once again, see artifactual hyperintensities at the edges when using the SVD basis, which are reduced when using the proposed basis (cf. magnifications in Fig. 4 k vs. l). The remaining qMT parameter maps for both participants can be found in Supporting Figs. S9-S10.
Four additional examples of qMT maps corresponding to different levels of motion ($25^{th}$, $50^{th}$, $75^{th}$, $85^{th}$ and $95^{th}$ percentile) can be found in Supporting Figs. S11–S20. With motions scores between the extreme cases shown in Fig. 5, the resulting parametric maps also show artifact levels in between these extremes. Overall, we observe that the proposed method either improves the image quality, or has similar image quality compared to the SVD basis.
To evaluate the performance of the motion correction across all 85 scans in our dataset, we calculated the standard deviation for the ROIs mentioned above. We analyzed them as a function of the respective pair-wise motion score (Fig. 6). Without motion correction, the standard deviation consistently increases with increasing motion, and the slope of a linear regression model differs from zero at a significance level of 0.01 for the majority of parameters and ROIs (see also Supporting Fig. S21). When performing motion correction, the motion-induced parameter variability is substantially reduced, with most of the regression slopes being non-significantly different from zero, which indicates effective motion correction with either basis.
Based on visual inspection of Fig. 6, the slope of the linear regression appears to be smaller when using the proposed basis in most ROIs and parameters, suggesting improved motion correction. To further evaluate this, we normalized the slope of each parameter and ROI by the respective intercept (no motion). We compared the three reconstructions, pooled over all qMT parameters (6) and ROIs (4) (Fig. 7). A paired t-test revealed a significant reduction in the normalized slope ($p<0.01$) when using the proposed basis compared to the SVD basis.
5 Discussion
We proposed enhancing the contrast-to-noise ratio between brain parenchyma and CSF by rotating the SVD basis. To this end, we used a generalized eigendecomposition, which is inspired by Region-Optimized Virtual Coils (ROVir).15 We demonstrated that the increased contrast overall provides improved accuracy in the motion estimates compared to an SVD basis, generally resulting in better image quality in the parametric maps.
To quantify the improvement in motion estimation accuracy provided by the proposed contrast-optimized basis, we performed simulations applying motion parameters to reference data with virtually no motion. We observed a decrease in the RMSE score when using the proposed contrast-optimized basis compared to the SVD basis in all four motion patterns.
@@ 5 Discussion
Figure 7: Analysis of the parameter estimates’ standard deviation in various ROIs as a function of the motion score (cf. Fig. 6). The histogram shows the slope of a linear regression, normalized by the respective intercept. This analysis pools all 6 qMT parameters and 4 ROIs. The proposed contrast-optimized basis results in smaller slopes compared to the SVD basis, indicating better motion correction.
@@
The proposed basis can be used as a one-to-one replacement for a traditional SVD basis. Therefore, both approaches have the same target applications, foremost transient-state quantitative MRI techniques, such as MR-Fingerprinting, where the same spin dynamics are repeated while filling the k-space. Furthermore, the proposed approach incurs no additional computational costs during reconstruction. Although we tested our method only on brain images and limited our investigation to rigid motion correction, the contrast-optimized basis can also be generated for other body parts with two distinct tissue types.
@@
Like with the original SVD approach, key factors for successful implementation include 3D imaging to mitigate through-slice motion and a k-space trajectory that provides adequate coverage in each repetition of the spin dynamics, facilitating a time-segmented reconstruction. In this study, we used a koosh-ball trajectory with golden-angle increments, which repeatedly samples the center of k-space. We paired it with a TV-regularized low-resolution reconstruction to mitigate undersampling artifacts.
@@
While the proposed motion correction can substantially reduce the artifacts in the parametric maps, Fig. 5 suggests that, in cases of severe motion, the image quality is still impaired despite motion correction with either basis. This is more evident in Fig. 5 and Supporting Fig. S10, which showed the largest amount of motion across all datasets (motion score of 10.53 mm): despite the considerable improvements obtained after motion correction (especially in the $R_{1}^{f}$ maps), the quantitative maps are still severely affected by motion artifacts. One explanation might be the inherently low temporal resolution of self-navigated motion correction. In our current implementation, each motion state is assigned every 4 s block, and improving the temporal resolution will be part of future work. Another source might be coil sensitivities, which are not adapted during the motion-corrected image reconstruction.
@@
The low temporal resolution entails the assumption that no motion occurs during a 4 s block. To address this limitation, we discard data before and after strong motion. However, in cases of continued strong motion (i.e., substantial variation of the motion states in each 4 s window), the trade-off between motion and undersampling artifacts imposes a ceiling on the motion correction performance.
@@ 6 Conclusions
We propose a contrast-optimized basis function for self-navigated motion correction in quantitative MR. We utilize the generalized eigendecomposition to enhance the contrast-to-noise ratio between brain tissues and CSF, which improves the accuracy of motion estimates and, ultimately, the image quality of the quantitative parametric maps. The proposed technique does not require any sequence modifications or additional scan time. Consequently, it can be seamlessly integrated into various quantitative MRI methods, e.g., inversion recovery or multi-echo spin echo, where signal variations over time can be effectively captured in a low-rank subspace.
@@
The reconstruction pipeline described in Sec. 3.2 is implemented in Julia v1.11, except for the volume registration (as part of the motion estimation) which is performed using SPM12 (Statistical Parametric Mapping) in MATLAB. The codes to generate the proposed contrast-optimized basis are available on https://github.com/EliMarchetto/ContrastOptimized_basis. Unfortunately, we cannot share the raw data, since facial recognition software could be used to identify the participants.
@@
Figure S1: In our acquisition scheme, we use six variable flip-angle patterns to encode six biophysical magnetization transfer (MT) parameters. Here we show the corresponding coefficient image derived using the SVD basis (a-f) and the proposed contrast-optimized basis (g-l). We observe considerable variability in the contrast between the six flip angle patterns, which can be attributed to a lack of contrast consideration in the optimization objective. In contrast, the proposed basis directly maximizes the contrast between tissues (in this case, brain parenchyma and CSF), resulting in less contrast variability. The images are reconstructed by aggregating all the radial spokes acquired during one 4 s RF cycle. More details on the reconstruction in Sec. 3.2.
@@
Figure S2: Distribution of the amount of data removal over the 86 datasets. The data removal is based on a motion score 28 and was calculated separately for the motion estimates derived using the SVD basis 14 and the proposed contrast-optimized basis.
Figure S3: RMSE scores for simulated motion, calculated by comparing motion estimates to the ground truth. For both the SVD basis and the proposed contrast-optimized basis, we tested ten different regularization parameters $\lambda$ to identify the optimal $\lambda$ (star markers) for each flip-angle pattern. The proposed basis yields lower minimum RMSE scores for most flip-angle patterns across three motion patterns.
Figure S4: Estimates of moderate simulated motion (motion score of 1.77 mm). The ground truth was obtained from an online database. The RMSE score, which quantifies deviations of the motion estimates from the ground truth (Sec. 3.3.3) is 0.38 mm when using the SVD basis and 0.35 mm when using the proposed basis. Downsampling the ground-truth motion to 4 s windows provides a lower bound of 0.3 mm for the RMSE score, given the low temporal resolution.
Figure S5: Parameter maps in the presence of moderate simulated motion, complementing the $m_{0}^{s}$ and $R_{x}$ maps in Fig. 3 of the main manuscript. The underlying data was corrupted with the pattern shown in Fig. 2 and Supporting Fig. S4. The RMSE for each parameter map, compared to the reference (no motion added), is shown in the top left of the respective map. The RMSE was calculated for the entire 3D brain volume, defined by the brain mask.
Figure S6: Estimates of strong simulated motion (motion score of 6.39 mm). The ground truth was obtained from an online database. The RMSE score, which quantifies deviations of the motion estimates from the ground truth (Sec. 3.3.3) is 1.33 mm when using the SVD basis and 1.26 mm when using the proposed basis. Downsampling the ground-truth motion to 4 s windows provides a lower bound of 1 mm for the RMSE score, given the low temporal resolution.
Figure S7: Parameter maps in the presence of strong simulated motion. The underlying data was corrupted with the pattern shown in Fig. S6. The RMSE for each parameter map, compared to the reference (no motion added), is shown in the top left of the respective map. The RMSE was calculated for the entire 3D brain volume, defined by the brain mask.
Figure S8: Estimates of inherent motion using the SVD basis 14 and the proposed contrast-optimized basis. We compare here a case with minimal ($0^{th}$ percentile motion score) and very strong ($100^{th}$ percentile motion score) unintended motion. The corresponding parametric maps can be found in Fig. 5 and Supporting Figs. S9 and S10.
Figure S9: Parameter maps in the presence of minimal inherent motion (0th percentile motion score), complementing the $m_{0}^{s}$ and $R_{x}$ maps in Fig. 5 of the main manuscript. In this case, all three reconstructions have virtually identical results.
Figure S10: Parameter maps in the presence of very strong inherent motion (100th percentile motion score), complementing the $m_{0}^{s}$ and $R_{x}$ maps in Fig. 5 of the main manuscript.
Figure S11: Estimates of inherent motion using the SVD basis 14 and the proposed contrast-optimized basis. We show here a case with a small amount of unintended motion (motion score 0.65 mm, $25^{th}$ percentile; no data removed). The corresponding parametric maps can be found in the Supporting Fig. S12.
Figure S12: Parameter maps in the presence of small inherent motion (25th percentile motion score). The corresponding motion parameters can be found in the Supporting Fig. S11).
Figure S13: Estimates of inherent motion using the SVD basis 14 and the proposed contrast-optimized basis. We show here a case with an average amount of unintended motion (motion score 0.91 mm, $50^{th}$ percentile; 1.11% of data removed). The corresponding parametric maps can be found in the Supporting Fig. S14.
Figure S14: Parameter maps in the presence of average inherent motion (50th percentile motion score). The corresponding motion parameters can be found in the Supporting Fig. S13).
Figure S15: Estimates of inherent motion using the SVD basis 14 and the proposed contrast-optimized basis. We show here a case with a substantial amount of unintended motion (motion score 1.28 mm, $75^{th}$ percentile; no data removed). The corresponding parametric maps can be found in the Supporting Fig. S16.
Figure S16: Parameter maps in the presence of substantial inherent motion (75th percentile motion score). The corresponding motion parameters can be found in the Supporting Fig. S15).
Figure S17: Estimates of inherent motion using the SVD basis 14 and the proposed contrast-optimized basis. We show here a case with a substantial amount of unintended motion (motion score 1.82 mm, $85^{th}$ percentile; 1.11% of data removed). The corresponding parametric maps can be found in the Supporting Fig. S18.
Figure S18: Parameter maps in the presence of substantial inherent motion (85th percentile motion score). The corresponding motion parameters can be found in the Supporting Fig. S17).
Figure S19: Estimates of inherent motion using the SVD basis 14 and the proposed contrast-optimized basis. We show here a case with an extreme amount of unintended motion (motion score 5.17 mm, $95^{th}$ percentile; 6.11% of data removed). The corresponding parametric maps can be found in the Supporting Fig. S20.
Figure S20: Parameter maps in the presence of extreme inherent motion (95th percentile motion score). The corresponding motion parameters can be found in the Supporting Fig. S19).
@@
Figure S21: Analysis of noise-like artifacts across all 85 datasets, complementing the analysis of $m_{0}^{s}$ and $R_{x}$ in Fig. 6. Each dot represents the standard deviation of the qMT parameter for one participant. The standard deviation was calculated for each motion correction version methods. Linear regression was performed to analyze the increase in the standard deviation with increasing motion. Black stars denote slopes that differ significantly from zero (* for p-value < 0.05, ** for p-value < 0.01). The slopes of this linear regression analysis are further analyzed in Fig. 7.
See pages - of SupportingInformation.pdf